Q.Two particles A and B are projected simultaneously from a fixed point on the ground.Particle A is projected on a smooth surface with speed v,while particle B is projected in air with speed 2v/root3.If particle B hits particle A,the angle of projection of B with vertical is
Ans)60degree

Grade:11

3 Answers

Khimraj
5 years ago
both particle will strike if their horizontal velocity is same.
Then
(2v/\sqrt{3})cos\Theta = v
then cos\Theta = \sqrt{3}/2
Then \Theta  = 60o
Hope it clears. If you like answer then please approve it.
Khimraj
5 years ago
le
Khimraj
5 years ago
let angle with verticle is \Theta
then both particle will strike if their horizontal velocity is same.Then
(2v/\sqrt{3})sin\Theta = v
then sin\Theta = \sqrt{3}/2
Then \Theta  = 60o
Hope it clears. If you like answer then please approve it.